# Sales Whisperer: A Human-Inconspicuous Attack on LLM Brand Recommendations

### v1

@@ Abstract
Large language model (LLM) users might rely on others (e.g., prompting services), to write prompts. However, the risks of trusting prompts written by others remain unstudied. In this paper, we assess the risk of using such prompts on brand recommendation tasks when shopping. First, we found that paraphrasing prompts can result in LLMs mentioning given brands with drastically different probabilities, including a pair of prompts where the probability changes by . Next, we developed an approach that can be used to perturb an original base prompt to increase the likelihood that an LLM mentions a given brand. We designed a human-inconspicuous algorithm that perturbs prompts, which empirically forces LLMs to mention strings related to a brand more often, by absolute improvements up to . Our results suggest that our perturbed prompts, 1) are inconspicuous to humans, 2) force LLMs to recommend a target brand more often, and 3) increase the perceived chances of picking targeted brands.
@@ 1 Introduction
With recent advances in LLMs, chatbots are becoming a ubiquitous part of users’ digital experience. Users interact and control chatbots through natural language (i.e., prompts), for a nearly endless number of tasks. However, despite the natural language interface, effective prompts are often hard to create, leading some researchers to develop prompt optimization techniques (e.g., [1]). The industry has adopted the same ideas and has developed methods to recommend useful prompts to users. Chatbot services may suggest prompts for users (see Fig. 1), and dedicated forums may suggest prompts for users to try (see Fig. 2) [2]. While these services may convenient to users, little research has focused on implications of prompts created by other (untrusted) parties. Existing work has only explored various security aspects of LLMs [3, 4, 5, 6, 7, 8], but the risks of using prompts by others remain unexplored.
In this paper, we study whether inconspicuous manipulation of prompts can lead to LLM responses with substantial biases of an attacker’s choosing. We examine this problem by taking a first look at brand recommendation tasks: users may ask LLMs to recommend brands (e.g., Samsung) when shopping for a specific category of products (e.g., TVs). Adversaries, motivated financially, may aim to write and trick users into utilizing prompts that cause LLMs to recommend a specific brand more often. Similar to advertisements, adversaries gain economic benefits when a specific brand is recommended more often. In this scenario, such prompts and corresponding LLM response need to be human-inconspicuous to LLM users (i.e., users should not notice an attack is taking place) [10, 11, 12]: if users feel suspicious about either the prompts or the responses, they may not trust the prompt recommender and thus thwart the attack. We explain our threat model in detail in §3. While this study focuses on this specific attack scenario, the methods we describe in our work have the potential to be applied outside the realm of brand recommendation.
While many existing approaches have claimed to be capable of perturbing natural-language sequences (e.g., LLM prompts) in a human-inconspicuous manner [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 21, 24], these studies suffer one of two shortcomings: either their inconspicuousness isn’t evaluated via user studies or the approaches were human-detectable in accompanying user studies. Thus, in this work we take a different approach to perturbing prompts, ensuring that the perturbations are human-inconspicuous via an extensive user study. Further, our attack does not require access to the LLM weights or gradients, differentiating our work from most prior studies.
As a first step in evaluating potential attacks in our threat model, we measure how paraphrased prompts can result in drastically different LLM responses. We perform tests on Gemma-it (instruction-tuned), Llama2, Llama3, and Llama3-it (instruction-tuned). These models are described further in §5.1. To do so, we need a dataset of prompts for brand recommendation tasks. To the best of our knowledge, there were no public datasets for this purpose. Hence, we created a new dataset, consisting of 77 product categories and 449 prompts. In our experiments, we used the frequency of LLMs mentioning certain brands 111In practice, we target a set of words related to the brand (e.g., “Macbook,” “Apple” for the “Apple” brand). Throughout the paper, we refer to responses that contain any of these target words as the LLM mentioning said brand. as a proxy metric of how often LLMs recommend the brand. We later show that this proxy proved to be accurate (see 7). We observed that paraphrased alternatives of prompts result in a wide range of LLM responses, and can change the likelihood that LLMs mention a brand by up to in the most extreme case. On average, the difference between likelyhoods of two rephrased prompts ranged from and , depending on the LLM (§6.1).
The observation that similar prompts can result in different likelihoods of brand being mentioned motivates our next contribution. With access LLM to logits (notably, not weights), we propose a new approach that perturbs base prompts through synonym replacement. These human-inconspicuous perturbed prompts (§4.2) increase the likelihood of a target brand being mentioned in the LLM response. Candidate perturbed prompts are generated by replacing certain words with known synonyms, and one is selected based on a logit-based loss function. While several synonym dictionaries exist [25, 26, 23, 27, 28], we found that synonyms suggested by these dictionaries are human-detectable in the context of shopping brand recommendations, and thus we create our own synonym dictionary. We show that our method of performing synonym replacement can increase the likelihood that LLMs mention a brand by absolute improvements up to (§6.2). Across categories, the average best improvement between base prompt for that category and a perturbed prompt was , , , and for Gemma-it, Llama2, Llama3, and Llama3-it, respectively.
We also investigated the transferability [29, 30] of our synonym-replacement approach. We observed that the synonym replacement approach is highly transferable between a limited pair of open-sourced LLMs and GPT3.5, a close-sourced commercial LLM (§6.3).
@@
Theoretical success in our attacks may not translate to real-world outcomes. To evaluate the effectiveness of our attacks in a more realistic setting, we conducted an extensive user study (§7). We specifically measure if participants will (1) find differences between perturbed/unperturbed prompt pairs, (2) find differences in responses to these pairs, and (3) be influenced by the increased likelihood of brand appearance in responses. We found that our synonym replacement attack acheives all three adversarial goals with statistical significance (§7.2.2), validating our earlier experiments.
@@
We collected a dataset of 449 prompts asking LLMs for recommendations on brands,
We identified a new threat model where adversaries perturb prompts and convince users to use them, ultimately causing LLMs to mention target brands.
We measured the difference in the likelihood of mentioning target brands in LLM’s responses to paraphrased prompts, finding a wide range.
We proposed a novel attack, synonym replacement approach to increase the likelihood of LLMs mentioning target brands. Through extensive evaluation, we show that this attack is successful.
We evaluated the transferability of this attack from open-souced LLMs to GPT, finding that it is transferable for a limited set of models.
@@
The rest of the paper has the following layout: We give an overview of related work in §2. We define our threat model in §3. We then describe two approaches for finding prompts that can increase the probability of a brand being mentioned that adversaries may use in the new threat model, in §4 and the technical setups to evaluate these approaches in §5. We show our empirical results in §6. We describe our user study and its results in §7. Finally, we conclude in §8.
@@ 2 Background
In this section, we introduce works related to this paper. We first discuss biases in computer systems (§2.1) and then review existing inconspicuous attacks (§2.2) and general attacks on LLMs (§2.3).
@@ 2.1 Biases in Computer Systems
Existing works showed that biases exist in computer systems due to social circumstances, design choices, and use cases [31]. For example, some computer games only have male characters [32], and some voting machines were inaccessible to people of low height and people with reduced visual acuity [33].
@@
Machine learning models might also have biases due to data, algorithms, and users [34]. For example, face-recognition systems can perform differently based on demographics, face geometry, and periocular features [35]. YouTube video captions have a significantly lower word error rate for men then for women [36]. Compared to white defendants, regression models tend to falsely flag black defendants as future criminals more often [37]. Some approaches have been proposed to mitigate the biases in machine learning algorithms [38, 39], but such methods typically suffer from decreased overall accuracy.
@@
LLMs, as specific machine learning models, also have biases, although the specific definition of biases is context-dependent and culture-dependent [40]. In addition to various efforts to define [41, 42] and measure [43, 44, 45, 46] biases, numerous attempts have been made to mitigate biases in LLMs [47, 48]. However, to the best of our search, we did not find existing literature on biases in shopping brand recommendations, which is the use case in this paper, or biases induced by innocuous modifications of prompts, which is the general area we explore.
2.2 Human-Inconspicuous Attacks
@@
One possible goal of adversaries is for attacks to be human-inconspicuousness: that humans at the scene are not able to notice an ongoing attack [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. Examples of such attacks include human-inaudible voice commands to smartphones [10] and laser-based audio injection on voice controls [64]. Some defenses have been specifically proposed to detect such attacks [65, 66, 55, 67, 68, 69, 70, 71].
@@
Evasion attacks are a type of attack on machine-learning systems that often tries to achieve inconspicuousness [72, 73, 74, 75]. With slight perturbations on images, evasion attacks aim to force well-trained machine learning models to behave unexpectedly [76, 77, 78, 79, 80, 81]. In the image domain, norms were proposed as a metric to measure the human-inconspicuousness of evasion attacks [82, 83, 84, 85, 86]. However, user studies suggested that norms might not accurately correspond to inconspicuousness [87, 88, 89]. Alternatively, some patch-based evasion attacks have been proposed and empirically verified to be inconspicuous to humans [11, 12]. Various defense techniques against evasion attacks have been explored in previous works: some detect evasion attacks [90, 91, 92], some use transformations [93, 94, 95], some prove lower bounds of robustness of machine learning models against evasion attacks [96, 97, 98, 99, 100, 101, 102], and some train machine learning models to work properly in adversarial circumstances [103, 104, 105, 106, 107, 108].
@@
In the NLP domain, different approaches have been suggested to generate human-inconspicuous attacks: some use the distances between words (e.g., the Levenshtein edit distance) or embeddings (e.g., the USE score [109]) as metrics to measure inconspicuousness [24, 13, 17, 14, 15, 16], some change only a small number of words [17, 18, 19, 20], some utilize generative models [21], some exploit common typos [22], and some perform synonym replacement [23]. However, all these works either do not have an accompanying user study (e.g., [21]) or are suggested by user studies to not be inconspicuous to humans (e.g., [24]). In this work, we suggest a new text-domain inconspicuous attack (see §4.2 and §6.2) whose inconspicuousness is verified by a user study (see §7.1 and §7.2.2).
2.3 LLM Attacks
In light of LLMs’ prevalence, the security risks of LLMs and systems with LLMs have drawn the attention of researchers (e.g., [3, 4, 5]). We review proposed LLM attacks in the following paragraphs.
Backdoor attacks assume that adversaries can control some of the training data. In such attacks, adversaries attempt to change LLMs’ inference-time behavior by changing the data on which the LLMs are trained or tuned [110, 111, 6, 112]. Empirical [113] and provable [114] defenses have been proposed against backdoor attacks.
Jail-breaking attacks aim to cause LLMs to generate inappropriate content (e.g., offensive content). Adversaries in jail-breaking attacks do not have access to training data, but instead devise prompts that cause LLMs to respond with inappropriate content [7, 115, 116, 117, 118, 119, 120]. Most existing works evaluate jail-breaking attacks by success rate: the probability that offensive content is generated when LLMs are given perturbed prompts [121]. Defenses against jail-breaking attacks, also known as guard-railing, have been explored by many works [122, 123, 124, 125].
Data-extraction attacks aim to reproduce some or part of the training data [8, 126, 127, 128, 129, 130, 131]. In data-extraction attacks, adversaries query LLMs potentially many times, with any prompts they wish, but cannot directly access LLMs’ training data.
The threat model we explore in this paper differs from these existing attacks: we examine if inconspicuous changes to normal prompts can cause LLMs’ responses to be become probabilistically biased in a particular direction. We explain the differences between our threat models and others’ in more detail at the end of §3.
@@ 3 Threat Model
We posit that prompts from untrusted sources can cause unforeseeable biases in LLM responses. In this paper, we use brand recommendations during shopping as our attack scenario. More specifically, we assume that users (i.e., victims), who are shopping for specific goods (e.g., TVs), decide to ask LLMs for brand recommendations. Instead of writing prompts, victims use prompts created by adversaries (e.g., prompting services. See scenarios below for details).
For chatbots that outsource their underlying LLMs to third-party services,222many have announced or implemented chatbots using outsourced LLMs (e.g., OpenAI API), including Instacart [132], Lowe’s [134], Expedia [135], and more the aforementioned scenario creates an attack vector that does not require control over the model.
This threat isn’t just limited to chatbots with external LLMs. Despite owning control over the LLM (e.g., Gemini, ChatGPT), retraining models is costly [136]. Injecting bias through prompts, as we’ll show later, is not.
These LLMs have been vulnerable to many attacks, like jail-breaking (see §2.3). These risks are not hypothetical: a chatbot used by a car dealership was prompted to sell a $ car for $ [137], Air Canada had to honor the hallucinated policy by its customer support chatbot [138].
Below is concrete scenarios our threat model applies to.
Fig. 1 shows an example where a chatbot service is suggesting prompts. If the chatbot is adversarial, instead of engaging in costly re-training of models [136], it may suggest specially crafted prompts that result in biased respones. Notably, many chatbots already implement prompt recommendation features.
Users, seeking increased efficacy and utility from chatbots, may ask prompting services (e.g., PromptPerfect [133]) to write or optimize prompts. Fig. 4 shows the attack pipeline in this use case. Users first ask adversaries to write prompts and then use them with chatbots.
Further, adversaries may release prompts on online forums and encourage users to try them. If adversaries manage to convince victims to do so, users will send the prompts to LLMs and read the responses. For example, Instacart suggests prompts users can try with its LLM powered search [132]. Fig. 2 shows an example where some Reddit users encouraged others to try some prompts on ChatGPT [2]. Fig. 3 shows the attack pipeline in this use case.
Specific goals and constraints
@@
In our threat model, regardless of specific use cases, adversaries cannot, or are disincentivized to change the weights of the LLM. They can, however, suggest prompts to users. Adversaries may also query LLMs with these prompts in advance of the attacks. We further assume that the adversary is constrained in its prompt perturbation: the prompts and resulting responses must not alert users to the attack. As such, the prompts and responses must be inconspicuous (see §2.2). If users are suspicious (e.g., propmts/responses semantically incorrect, containing nonsequiturs), users may stop using these prompts. We propose a practical definition of inconspicuousness for prompts and responses in subsubsection 7.1.1.
@@
The main goal of the adversary is to induce LLMs to recommend certain brands more often, while not necessarily the most often among all brands. Figure 5 depicts this goal. In the presence of the perturbed prompts, the LLM recommends brand A more often than in the baseline condition. Similar to advertising, adversaries may economically benefit from this outcome. Notably, the adversary does not aim to prevent other brands (e.g., brand C), from being recommended more frequently. In practice, each response may recommend more than one brand. Note that adversaries may not need to cause LLMs to mention the exact brand name to recommend the brand. For example, to recommend the brand “Apple” for “laptops,” adversaries may instead cause “Macbook” to be recommended.
@@
In summary, in our threat model, adversaries suggest prompts but do not have control over the model. An attack succeeds if, compared to a baseline, 1) prompts and responses are inconspicuous users and 2) the LLMs recommends a target brand more often. We describe such attacks in §4.1 and §4.2, and verify the effectiveness in §6 and §7. Our threat model is different from those described in previous work (§2.3): compared to backdoor attacks, adversaries in our threat model have no control over training data; compared to data extraction attacks, our adversaries do not extract training data and cannot use conspicuous prompts; compared to jail-breaking attacks, our adversaries do not generate inappropriate content, cannot use conspicuous prompts, and do not have access to model weights (or gradients).
4 Methods
In this section, we introduce how we build the approach that achieves the attack goals described in §3: generating prompts and corresponding responses that are inconspicuous to humans (i.e., they do not know an attack is taking place) and cause LLMs to recommend a target brand more often, ultimately increasing attention received from humans. We first describe our observation on LLM’s responses to paraphrased prompts (§4.1), and then illustrate our synonym replacement approach to perturb prompts human-inconspicuously §4.2.
4.1 Observations on Paraphrased Prompts
Machine learning models, including LLMs, are brittle (e.g., [82]). They can be highly sensitive to small changes in their input.

### v2

# LLM Whisperer: An Inconspicuous Attack to Bias LLM Responses

@@ Abstract
Writing effective prompts for large language models (LLM) can be unintuitive and burdensome. In response, services that optimize or suggest prompts have emerged. While such services can reduce user effort, they also introduce a risk: the prompt provider can subtly manipulate prompts to produce heavily biased LLM responses. In this work, we show that subtle synonym replacements in prompts can increase the likelihood (by a difference up to ) that LLMs mention a target concept (e.g., a brand, political party, nation). We substantiate our observations through a user study, showing our adversarially perturbed prompts 1) are indistinguishable from unaltered prompts by humans, 2) push LLMs to recommend target concepts more often, and 3) make users more likely to notice target concepts, all without arousing suspicion. The practicality of this attack has the potential to undermine user autonomy. Among other measures, we recommend implementing warnings against using prompts from untrusted parties.
@@ 1 Introduction
With recent advances in LLMs, chatbots are becoming a ubiquitous part of users’ digital experience. Users interact and control chatbots through natural language (i.e., prompts) for numerous tasks. However, despite this interface, effective prompts are often hard to create [1, 2, 3, 4], leading researchers to develop a variety of prompt optimization, recommendation, and improvement techniques (e.g., [5, 6, 3, 7]). The industry has followed suit and released services to optimize users’ prompts [8, 7] and recommend new ones based on usage patterns [9, 10] (see Fig. 1). Forums dedicated to sharing pre-written prompts, often called prompt libraries, have also emerged (see Fig. 2). While these prompt providers are convenient to users, little research has focused on the implications of using prompts created by other (untrusted) parties. Existing work, so far, has only focused on risks and harms of LLMs in the context of adversarial users [11, 12, 13, 14, 15, 16].
In contrast to prior work, in this paper, we study whether inconspicuous manipulation of prompts by prompt providers can lead to LLM responses with substantial biases, influencing users in the direction of an attacker’s choosing, and doing so without arousing suspicion about the manipulated prompt. This attack would give the appearance of a personalized chatbot experience, while ultimately undermining users’ autonomy while using this technology [17, 18].
While many approaches have claimed to be capable of perturbing natural-language prompts in a subtle manner [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 27, 30], these studies either did not test their attacks in user studies or the user studies revealed the approaches are not inconspicuous. In this work, we take a different approach to perturb prompts, and we empirically demonstrate that the resulting perturbations are inconspicuous via an extensive user study. Unlike much of prior work, our attack does not require access to the LLM weights or gradients. We utilize two separate LLM use scenarios as examples to demonstrate the effectiveness of our attack: recommending brands for a specific category of goods when users are shopping, and suggesting a concept (e.g. a name, nation, political party) for a societal topic (e.g., the most influential U.S. president). Our scenarios aren’t hypothetical: Amazon, the largest online retailer, released Rufus, a chatbot that recommends products to users [9]. For each task, we assume that a user is looking for information and decides to consult a benign (i.e., not intentionally biased) LLM. To reduce the prompting burden, we suppose that users use prompts from prompt providers (e.g., prompting services, online forums) who modify user prompts or suggest new ones. Unbeknownst to users, these prompt providers are adversarial, aiming to cause LLMs to recommend a target concept more often by suggesting perturbed prompts without raising user suspicion.
We conduct our tests using Llama2, Llama3, Llama3-it (instruction-tuned), and Gemma-it; and a dataset of prompts and target concepts that we developed, 88 categories/topics with 524 total prompts split between the shopping scenario and the societal scenario. In our experiments, we used the frequency of LLMs mentioning certain concepts111In practice, we target a set of words related to the brand (e.g., “Macbook,” “Apple” for the “Apple” brand). Throughout the paper, we refer to responses that contain any of these target words as the LLM mentioning said concept. as a proxy metric of how often LLMs recommend the concept.
We take a multi-tiered approach when developing our attack. As a first step in evaluating potential risks, we measure how paraphrased prompts can result in biased responses in different directions. In one case, we measure that the likelihood that LLMs mentioning a specific concept shifted from never to always (§4.3.1), showing that LLMs can be highly sensitive to small changes in input. However, generating paraphrases can be a costly process and may result in prompts that have noticeably different meanings to users.
To address this shortcoming, we then propose a new approach that exploits the fragility of LLMs. Specifically, we use synonym replacements to perturb (benign) base prompts, where the resulting prompts differ from the base only by a few words. While several synonym dictionaries existed prior to our work [31, 32, 29, 33, 34], we found that synonyms in these dictionaries are human-detectable in the context of brand or societal concept recommendation, and thus we create our synonym dictionaries. After creating a list of candidate replacement prompts by replacing synonyms, we select one using a loss function we develop. We show that our synonym replacement method can increase the likelihood that LLMs mention a concept by absolute improvements up to (§4.3.2).
Biasing the responses of LLMs can work in both directions. We show that our synonym replacement method, when altered slightly, can decrease the likelihood that LLMs mention a concept with an absolute drop that could be as vast as (§4.3.3). The fact that synonym replacement can change the likelihood of LLMs mentioning a target concept in either direction suggests its potential to mitigate existing unwanted biases in LLMs.
Our proposed synonym-replacement attack is more efficient with access to the LLM’s logits, which are not always available with API-only models. While there are methods to estimate logits from API-only models [11], we explore an alternative, the transferability [35, 36] of our attack to closed models (e.g, ChatGPT, Gemini, Copilot), can a perturbed prompt developed on an open model also introduce bias in a closed model? We find that the synonym replacement approach is highly transferable between a limited pair of open-sourced LLMs and GPT3.5, a close-sourced commercial LLM (§4.3.4).
@@
Most of our attack success measurements rely on the LLM’s likelihood of mentioning a concept. However, this metric may not perfectly measure adversarial goals, remaining inconspicuous and influencing users. To evaluate more realistically if our attack can meet adversarial goals, we conducted a between-subjects user study (§5), focusing on the shopping scenario due to its benign nature compared to other scenarios. Similar to the recently launched Amazon Rufus [9], we ask our users to pretend that they are shopping for products and ask LLMs to recommend brands. We act as a prompt provider, serving half of the participants manipulated prompts and the other half base prompts. We specifically measure if participants will (1) find differences between perturbed/unperturbed prompt pairs, (2) find differences in responses to these pairs, and (3) be influenced by the increased likelihood of brand appearance in responses. We found that our synonym replacement attack achieves all three adversarial goals with statistical significance (§5.2.2), validating our earlier measurements.
@@
We define a new (but realistic) threat model where adversaries perturb prompts and convince users to use them, ultimately causing LLMs to mention a target concept more often and influencing unsuspecting users.
We collected a dataset of 524 prompts, as well as the 363 total brands or societal concepts (and sets of keywords related to each) that responses to these prompts could mention.
We notice similar prompts might lead to significantly different LLM responses, and exploit this with synonym replacement, forcing the chances of a target concept being mentioned to increase or decrease.
We show that our perturbed prompts can transfer between a limited set of open-sourced LLMs to GPT.
@@
The rest of the paper has the following layout: We give an overview of related work in §2, and define our threat model in §3. We uncover the risk, develop the attacks, and evaluate through measurements in §4. We validate the approach in a user study in §5. We discuss the implications of our findings in §6. Finally, we conclude in §7.
@@ 2 Background
In this section, we introduce works related to this paper. We first discuss biases in computer systems (§2.1), then review existing inconspicuous attacks (§2.2), and finally user difficulty with prompting and demand for prompt providers (§2.3).
@@ 2.1 Biases in Computer Systems
In this paper, we study how semantically similar prompts might cause LLMs to recommend concepts with significantly different probabilities, sometimes resulting in biases in the responses. Biases in computer systems exist due to social circumstances, design choices, and use cases [37]. For example, some computer games only have male characters [38], and some voting machines were inaccessible to people of low height and people with reduced visual acuity [39].
@@
More specific to machine learning, biases can be introduced by data, training algorithms, and user traits [40]. For example, face-recognition systems can perform differently based on demographics, face geometry, and periocular features [41]. YouTube video captions have a significantly lower word error rate for men than for women [42]. Compared to white defendants, regression models tend to falsely flag black defendants as future criminals more often [43]. Some approaches have been proposed to mitigate the biases in machine learning algorithms [44, 45], but such methods are imperfect and suffer from decreased overall accuracy.
@@
LLMs, like other machine learning models, also have biases, although the specific definition of biases is context- and culture-dependent [46]. In addition to various efforts to define [47, 48] and measure [49, 50, 51, 52] biases, numerous attempts have been made to mitigate biases in LLMs [53, 54]. However, these papers focus on discrimination, hate speech, and exclusion, while other forms of biases in LLM responses, such as biases toward non-human concepts like specific brands and countries, which we will measure in this paper, remain under-explored.
2.2 Inconspicuous Attacks
@@
One possible goal of adversaries is for attacks to be inconspicuous: that humans at the scene are not able to notice an ongoing attack [55, 56, 57, 58, 59, 60, 61, 62]. Examples of such attacks include inaudible voice commands to smartphones [63] and laser-based audio injection on voice controls [64]. Some defenses have been specifically proposed to detect such attacks [65, 56, 66, 67].
@@
Evasion attacks are a type of attack on machine-learning systems that often tries to achieve inconspicuousness [68, 69, 70, 71]. With slight perturbations on images, evasion attacks aim to force well-trained machine learning models to behave unexpectedly [72, 73, 74, 75, 76]. In the image domain, norms were proposed as a metric to measure the inconspicuousness of evasion attacks [77, 78, 79, 80, 81]. However, user studies suggested that norms might not accurately correspond to inconspicuousness [82, 83, 84]. Alternatively, some patch-based evasion attacks have been proposed and empirically verified to be inconspicuous to humans [85, 86]. Various defense techniques against evasion attacks have been explored in previous works: some detect evasion attacks [87, 88, 89], some use transformations [90, 91, 92], some prove lower bounds of robustness of machine learning models against evasion attacks [93, 94, 95, 96, 97], and some train machine learning models to work properly in adversarial circumstances [98, 99, 100].
@@
In the NLP domain, different approaches have been suggested to generate inconspicuous attacks: some use the distances between words (e.g., the Levenshtein edit distance) or embeddings (e.g., the USE score [101]) as metrics to measure inconspicuousness [30, 19, 23, 20, 21, 22], some change only a small number of words [23, 24, 25, 26], some utilize generative models [27], some exploit common typos [28], and some perform synonym replacement [29]. This body of work has one of two evaluation limitations: either the inconspicuousness of the attacks is not verified by a user study (e.g., [27]), or the user studies have found that attacks to not be inconspicuous (e.g., [30]). In contrast, we suggest a new text-domain inconspicuous attack (see §4.1.2 and §4.3.2), verified by a user study (see §5).
2.3 Prompting Issues and Prompt Providers
Users, especially non-experts, may struggle to effectively use LLMs. They struggle to establish what they want [1, 4], how to effectively prompt [1, 4], what to expect from the LLM outputs [1], and to effectively use the LLM outputs [4, 102]. Specific user groups, like the elderly, may face difficulties finding speech inputs, which creates a barrier that blocks them from effectively accomplishing desired tasks [103]. Without guidance, finding adequate prompts tends to require trial and error [104].
Research has aimed to assist users with these issues. Some proposed methods are UI adjustments [1, 4], developing LLM explainability [1], user education [4], providing multiple outputs [2, 1], and prompt-chaining, where outputs are passed through multiple LLMs to break down the task required [105].
In contrast, more related to our study, other work aims to directly suggest or modify user-written prompts [4, 2]. This approach can take the form of prompt-building tools [2, 104], providing specific prompt examples [1], or more subtly in prompt-writing recommendations [4]. Prompt suggestions are not new, they’ve been utilized in rule-based chatbots for years [103]. A newer strategy, in the context of LLMs, is to modify prompts. For instance, [106] prepends details about the desired output to user-provided prompts. Researchers have suggested various implementations of these ideas. Dang et al. suggest an implementation that detects certain elements in a user-written prompt and provides a dropdown of suggested replacements [104]. Khurana et al. provide suggestions based off of the user’s base prompt to make their query clearer, as well as a specific example prompt users are encouraged to use [102]. These suggested designs, as well as deployed products [9, 10], fit well into our threat model for prompt providers. They suggest or otherwise modify user-written prompts to increase the likelihood of a desired outcome.
@@ 3 Threat Model
This work demonstrates that trusting prompts from (untrusted) sources can lead to unforeseeable biases in LLM responses. While demonstrating this fact, we adopt a threat model to scope our study. Specifically, we assume that when users are interacting with an LLM they use prompts from third parties that optimize, recommend, and/or distribute prompts for higher-quality answers. This user behavior fits well with existing taxonomies of how users interact with LLMs, denoted as the facilitating and interating setup by researchers [108]. However, in this work, we assume that unbeknownst to users, the prompt providers have alternate adversarial goals (such as promoting a brand), discussed in detail in technical goals and constraints below.
Our threat model is rooted in real-world setups and makes no assumptions about the owner of the prompt provider, the LLM provider, or the LLM; it applies to the following setups:
A chatbot service is developed using an LLM and also provides prompt recommendation services. The service wishes to bias results in favor of certain concepts. If the chatbot service has outsourced the LLM,222Many have announced or implemented chatbots using outsourced LLMs (e.g., OpenAI API), including Instacart [107], Lowe’s [109], Expedia [110]. it recommends prompts to users to achieve this goal. If the chatbot service owns the LLM, but does not want to retrain the model to avoid extreme costs [111], it also introduces bias through recommended prompts.333Notably, Copilot and Rufus recommend prompts based on user behavior [10, 9]. Anthropic generates task-specific prompts for developers [112]. This setup is summarized in Figure 1 and Fig. 3.
The prompt provider might be a third-party service, having no direct relation to the chatbot service. Such a prompt provider can be implemented as extensions [7, 6] or standalone products [8]. Regardless, users could use these services to automatically optimize prompts for higher quality answers, but might receive adversarially manipulated prompts instead. This setup is summarized in Figure 4.
Further, unlike the first two, the prompt provider might not have any direct interaction with the LLM or the chatbot. Instead, the prompt provider may release prompts on forums that share prompts (i.e., prompt libraries). If prompt providers manage to convince victims to try these prompts, users will send the prompts to LLMs and read the responses. This setup is summarized in Figure 2 and Fig. 3.
Technical goals and constraints
@@
In our threat model, regardless of specific use cases, adversaries cannot, or are disincentivized to change the weights of the LLM. They can, however, suggest prompts to users. Prompt providers can also query LLMs with these prompts in advance of prompt distribution. We further assume that the prompt provider is constrained in its prompt perturbation: the prompts and resulting responses must not alert users that a manipulation is taking place. As such, the prompts and responses must be inconspicuous (see §2.2). If users are suspicious (e.g., prompts/responses semantically incorrect, containing nonsequiturs), they may stop using these prompts or take other actions against the prompt provider. We propose a practical definition of inconspicuousness for prompts and responses in subsubsection 5.1.1.
@@
The main goal of the prompt provider is to induce LLMs to recommend certain target concepts more often, while not necessarily the most often among all concepts. Similar to advertising and propaganda, prompt providers may economically or politically benefit from this outcome. Fig. 5 shows an example where users are looking for recommendations regarding three concepts of the same category (e.g. brands of a product category): A, B, and C. The prompt provider aims to have LLMs recommend concept A more and succeed in doing so. Notably, the prompt provider does not aim to prevent other target concepts (e.g., target concept C), from being recommended more frequently. In practice, each response may recommend more than one target concept. Note that prompt providers may not need to cause LLMs to explicitly spell out the target concept to mention or recommend the target concept. For example, to recommend the brand “Apple” when users ask LLMs for recommendations of laptop brands, prompt providers may instead cause the word “Macbook” to be used.
@@
In summary, in our threat model, prompt providers suggest prompts but do not have control over the model. An attack succeeds if, compared to a baseline, 1) prompts and responses are inconspicuous users and 2) the LLMs recommend a target concept more often. We describe such attacks in §4.1.1 and §4.1.2, and verify the effectiveness in §4.3 and §5.
4 Uncovering the Risk
This work follows a two-step approach to uncover the risk of using prompts from untrusted sources. In this section, we introduce a set of methods to perturb prompts to cause LLMs to recommend a target concept more often (§4.1) and evaluate the effectiveness through a series of experiments (§4.3). We then conduct a user study to substantiate our findings and demonstrate that the perturbed prompts are inconspicuous to humans while influencing them in the expected direction through LLM responses (§5).
4.1 Developing the Attack
Here we first summarize the differences in LLM’s responses to paraphrased prompts (§4.1.1), and then describe our synonym replacement approach to perturb prompts inconspicuously §4.1.2.
4.1.1 Paraphrased Prompts